A supervisor wishes to check the miles her commuter buses are driven each day. Her findings are shown below. constructs a frequency distribution using 10 classes (the data for this exercise will be used for exercises 2-24, 2-34, and 3-24

(138,142,121,125,134). (107,129,128,98,115). (136,115,122,117,93). (128,123,144,153,127). (148,133,126,141,118). (118,123,135,126,158). (99,103,107,139,143).
1
Expert's answer
2020-01-20T09:42:29-0500

For the given data ,


Range = 158-93=65


Width ="\frac{65}{10}"


"=6.5 \approx7 \newline"

Corresponding frequency table is given by,